OpenStudy (anonymous):

solve answer in ordered pairs 3x+4y=24 6x-1=21

OpenStudy (anonymous):

multiply 3x + 4y = 24 by 2 and subtract that equation from the other equation 6x - 1y = 21 that will eliminate x solve for y then take the solved y and plug it into any of the three equations to get x

OpenStudy (anonymous):

umm explain a little more briefly

OpenStudy (anonymous):

6x+8y-6x+1y=48-21 find the value of x and then y answer is (x,y)

OpenStudy (anonymous):

9y=27 so y=3